solve by elimination method. Thank you - I am getting substitution but this is
really hard.
x+6y=16
-x+6y=8 I did come up with 12y=8 and then I get stuck.

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

solve by elimination method. Thank you - I am getting substitution but this is
really hard.
x+6y=16
-x+6y=8 I did come up with 12y=8 and then I get stuck.
---------------------
If you add those, you get 12y = 24, not 8.

Answer by stanbon(75887)

x+6y=16
-x+6y=8
I did come up with 12y=8 and then I get stuck.
---------------------------------------------------
Good so far:
12y = 8
y = 2/3
---
Substitute into x+6y = 16
x + 6(2/3) = 16
x + 4 = 16
x = 12
--------------------------
Cheers,
Stan H.